E) expected output is less than rated capacity 9. Which of the A) ap levels service rule in its computer labs to provide capacity of the system is 1900 students per E) early bird specials in restaurants 10. The Academic Computing Center has five trainers avai training sessions to students. Assume that the design of students who 90% of design capacity. If the number semester and that effective capacity equals actually got their orientation session is 1500, what is the ef of A) 1350 students B) 1710 students 78.9% 87.7% )90% time? tions that have process times of 5, 9, 4,9, and 8. What is the bottleneck A) 4 B) 9 C) 18 D) 35 E) 7 12. A work system ha s five stations that have process times of 5, 9, 4,9, a nd 8. What is the throughput time of the system? A) 4 B) 9 C) 18 D) 35
Answer #1

9. C Changes in staffing levels

To manage capacity, changes in staffing levels need to be done in order to match varying capacity.

10. D 87.7%

1900*90% = 1710

efficiency = 1500/1710*100 = 97.72%

11. B 9

Bottleneck time is the longest time.

12. D 35

5+9+4+9+8 =

It is the total time required for a product to pass the production process.